# Normalized solutions of Nehari-Pankov type to indefinite variational problems

### v1

@@ 5.2. Weyl’s law, large spectral gaps and large $L^{\infty}$ norms
$\lambda_{n}(u)=\inf_{\stackrel{{\scriptstyle E^{\prime}\subset E}}{{\dim E^{\prime}\geq n}}}\sup_{v\in E^{\prime}\setminus\{0\}}\frac{1}{\|v\|_{H}^{2}}\Bigl(\langle A^{1/2}v,A^{1/2}v\rangle-\int_{\Omega}V_{u}v^{2}\,\mathrm{d}x\Bigr).$ (48)
@@ 5.3. ODE arguments and the $L^{\infty}$ to $L^{2}$ estimate
$u^{\prime\prime}=\lambda u+f(u)$ (49)
@@
$\frac{\tau_{\lambda,M}}{8}\min(m_{\lambda,M},M)^{2}\leq\int_{0}^{\tau_{\lambda,M}}|u_{\lambda,M}(x)|^{2}\,\mathrm{d}x\leq\tau_{\lambda,M}\max(m_{\lambda,M},M)^{2}.$ (50)
@@
$\tau_{\lambda,M}=2\int_{-m_{\lambda,M}}^{M}\frac{\,\mathrm{d}u}{\sqrt{2(H_{M}-V_{\lambda}(u))}}.$ (51)
@@
$\min(m_{\lambda,M},M)\geq\frac{1}{\kappa_{0}}\max(m_{\lambda,M},M)\geq\frac{1}{\kappa_{0}}R$ (52)
@@ 6. The nonlinear biharmonic equation on the $2$-torus
$r_{2}(j)\leq C_{\varepsilon}j^{\varepsilon}\quad\text{for every $j\in\mathbb{Z}^{>0}$.}$ (53)
@@
$d(n_{\ell},\lambda(\ell))\leq C_{\tau}(\lambda_{n_{\ell}+1}-\lambda(\ell))^{-\tau}.$ (55)
@@
$\sup\Bigl\{\|u\|_{L^{2}}\bigm|\text{$u\in Q_{\lambda}$ for some $\lambda\in(\lambda_{n_{\ell}},\lambda_{n_{\ell}+1})$}\Bigr\}\geq\mu+1\qquad\text{for $\ell\geq\ell_{\mu}$.}$ (56)
@@
Combining the estimate (56) with Theorem 1.7 yields the existence of infinitely many normalized solutions of the equation (8) with mass $\mu$ corresponding to an infinite set of parameters $\lambda$ located in the disjoint intervals $(\lambda_{n_{\ell}},\lambda_{n_{\ell}+1})$, $\ell\geq\ell_{\mu}$. The proof of Theorem 1.4 is thus finished.

### v2

@@ 5.2. Weyl’s law, large spectral gaps and large $L^{\infty}$ norms
(50) $\lambda_{n}(u)=\inf_{\stackrel{{\scriptstyle E^{\prime}\subset E}}{{\dim E^{\prime}\geq n}}}\sup_{v\in E^{\prime}\setminus\{0\}}\frac{1}{\|v\|_{H}^{2}}\Bigl(\langle A^{1/2}v,A^{1/2}v\rangle-\int_{\Omega}V_{u}v^{2}\,\mathrm{d}x\Bigr).$
@@ 5.3. ODE arguments and the $L^{\infty}$ to $L^{2}$ estimate
(51) $u^{\prime\prime}=\lambda u+f(u)$
@@
(52) $\frac{\tau_{\lambda,M}}{8}\min(m_{\lambda,M},M)^{2}\leq\int_{0}^{\tau_{\lambda,M}}|u_{\lambda,M}(x)|^{2}\,\mathrm{d}x\leq\tau_{\lambda,M}\max(m_{\lambda,M},M)^{2}.$
@@
(53) $\tau_{\lambda,M}=2\int_{-m_{\lambda,M}}^{M}\frac{\,\mathrm{d}u}{\sqrt{2(H_{M}-V_{\lambda}(u))}}.$
@@
(54) $\min(m_{\lambda,M},M)\geq\frac{1}{\kappa_{0}}\max(m_{\lambda,M},M)\geq\frac{1}{\kappa_{0}}R$
@@ 6. The nonlinear biharmonic equation on the $2$-torus
(55) $r_{2}(j)\leq C_{\varepsilon}j^{\varepsilon}\quad\text{for every $j\in\mathbb{Z}^{>0}$.}$
@@
(57) $d(n_{\ell},\lambda(\ell))\leq C_{\tau}(\lambda_{n_{\ell}+1}-\lambda(\ell))^{-\tau}.$
@@
(58) $\sup\Bigl\{\|u\|_{L^{2}}\bigm|\text{$u\in Q_{\lambda}$ for some $\lambda\in(\lambda_{n_{\ell}},\lambda_{n_{\ell}+1})$}\Bigr\}\geq\mu+1\qquad\text{for $\ell\geq\ell_{\mu}$.}$
@@
Combining the estimate (58) with Theorem 1.8 yields the existence of infinitely many normalized solutions of the equation (8) with mass $\mu$ corresponding to an infinite set of parameters $\lambda$ located in the disjoint intervals $(\lambda_{n_{\ell}},\lambda_{n_{\ell}+1})$, $\ell\geq\ell_{\mu}$. Moreover, all these solutions are sign changing. Indeed, this follows as in Section 5.4 (or by simply integrating (8) over $\Omega=S^{1}\times S^{1}$), since the eigenspace of $A=\Delta^{2}$ associated with $\lambda_{1}=0$ is spanned by constant functions (which therefore do not change sign). The proof of Theorem 1.4 is thus finished.